OpenStudy (anonymous):

A TV remote has keys for channels 0-9. You select one key at random. What is the probability that you select a channel that is an odd number?

OpenStudy (anonymous):

P = (number of odd channels from 0-9) / total number of channels P = 5/10=0.5

OpenStudy (anonymous):

so does that mean 1 in 5?

OpenStudy (anonymous):

no 1 in 2

OpenStudy (anonymous):

so it's 1/2

OpenStudy (anonymous):

yes. there are 5 odd channels in a total of 10 channels so, the odds of randomly selecting an odd channel is 5/10=1/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yer welcome